if a man drives 70 miles per hour how far will he travel in 3
hours? in 4 hours? and in n hours

Answer by jim_thompson5910(35256)

We're using D = r*t each time. Each time, r = 70 while t will change


In 3 hours, he will travel 70*3 = 210 miles

In 4 hours, he will travel 70*4 = 280 miles

In n hours, he will travel 70*n = 70n miles